OpenStudy (anonymous):

There are 36 apples in a bag. If Yuri takes out 6 of the apples, in simplest form, what fraction of the apples did he take out of the bag?

OpenStudy (anonymous):

1/6

OpenStudy (anonymous):

Well he took 6 of the 36 apples, which is 6 / 36. 36 = 6 * 6 This means that we can divide them by 6 6 / 6 = 1 36 / 6 = 6 6/36 = 1/6

OpenStudy (anonymous):

We want to divide them by their GCD, http://www.wolframalpha.com/input/?i=GCD%2836%2C6%29